# Fixed Points of the Josephus Function via Fractional Base Expansions

### v2

@@ Abstract
In this paper, we investigate some interesting properties of the fixed point sequence of the Josephus function $J_{3}$. First, we establish a connection between this sequence and the Chinese Remainder Theorem. Next, we identify a clear numerical pattern for the digits of two consecutive fixed points when they are written in a non-standard fractional number system in base $3/2$. This result allows us to derive a recursive procedure for determining the digits of their base $3/2$ expansions.
@@ 1 Introduction
This work discusses some interesting properties of the Josephus function $J_{3}$, which arises from the classical Josephus problem, a well-known mathematical riddle originating from a narrative by the Jewish-Roman historian Flavius Josephus (see [4, Book 3, Chapter 8, Part 7]). The problem is described as follows: Given a group of $n$ people seated at a round table, numbered consecutively clockwise from $1$ to $n$, begin counting from position $1$, skip two people, and eliminate the third. Continue this process with the remaining participants, always starting with the next person, counting two, and removing the third, proceeding clockwise until only one person remains. This person is called the survivor.
A natural question is where one should initially sit to avoid elimination. Let $J_{3}(n)$ denote the position of the survivor. Computing $J_{3}(n)$ directly for large $n$ can be costly, so it is useful to identify features of $J_{3}$ that can be exploited; see [3, 2].
@@
Definition (High extremal points).
@@
Next, we present the main result of this paper, organized into three subsections. We begin by motivating the main goal of the manuscript by recalling the well-known explicit formula for the fixed points of $J_{2}$. Subsection 2.1 connects the search for fixed points of $J_{3}$ with the Chinese Remainder Theorem, representing each fixed point as a solution of a system of congruences with pairwise coprime moduli. Subsection 2.2 introduces a fractional number system in base $3/2$, following [1, 7], and studies its relevance to fixed points. Finally, Subsection 2.3 uses this system to recursively determine the digits in the base $3/2$ expansion of each fixed point of $J_{3}$.
@@ 2 Main Results and Motivations
In this paper, we aim to derive a recursive formula for the digits of consecutive fixed points for $J_{3}$ in a suitable base. Such a formula was obtained in full for the special case of $J_{2}$ in [6, 5]. In this case, the fixed points are given by the explicit formula $n_{p}^{(\ell)}=2^{\ell}-1$, for all $\ell\in\mathbb{N}$. Their expansions in base $2$ exhibit a particularly simple recursive pattern. Specifically, if
@@ 2.1 Fixed points of $J_{3}$ and the Chinese Remainder Theorem
which tells us that $x_{1}=n_{p}^{(\ell+1)}$ and $x_{2}=n_{p}^{(\ell)}$ are each a solution to the following system of linear congruences:
@@
where $p:=\overline{m}_{\ell}$ and $q:=\overline{m}_{\ell+1}$. Note that when $pq=0$, our argument does not apply in this case. From this point on, we assume that $p,q\geq 1$. Independently, one can prove by induction on $p$ and $q$ that
@@
where $\gcd$ is the greatest common divisor. In this way, we uncover an interesting connection between the sequence $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ and the CRT. Moreover, since the explicit expression for $z$ depends on a solution to Bézout’s identity, it is natural to examine how $1$ can be expressed as a linear combination of $3^{p}$ and $2^{q}$.
@@ 1 $1=3^{p}(1)+2^{1}\left(\tfrac{1}{2}(1-3^{p})\right),\quad\text{if }p\geq 1$
Table 2 provides Bézout’s linear combinations for all $p\in\mathbb{N}$ and for $q=1,2,3,4,5$ in a compact form. We invite the interested reader to extend Table 2 to other values of $q$.
Remark.
@@
The first example corresponds to the following values $n_{p}^{(17)}=3986218$ and $n_{p}^{(18)}=102162424$ given in [3, Table 1] or Table 2, where $p=7$ (i.e., $\overline{m}_{17}$) and $q=1$ (i.e., $\overline{m}_{18}$), respectively. According to (5) and Table 2, $a_{1}=1093$, $a_{2}=0$, and $x=1$, y=1093y=-1093. Therefore,
@@
Thus, $n_{p}^{(18)}=102162424$ must be equivalent to 3280 modulo 4374, which is indeed true since (1021624243280)/4374=23356(102162424-3280)/4374=23356.
Another example is obtained when $n_{p}^{(13)}=46084$ and $n_{p}^{(14)}=103690$; see Table 3 below. In this case, $p=1$ (i.e., $\overline{m}_{13}=1$) and $q=5$ (i.e., $\overline{m}_{14}=5$). Here, $a_{1}=1$, $a_{2}=10$ in (5), and $x=11$, y=1y=-1 for the Bézout identity (6) since
@@ 2.2 Expansions in base $3/2$
which provides motivation for studying representations of natural numbers in base $3/2$. Inspired by (7) and the basic conceptual framework presented in the introduction to [7, Section 3], we move from base 3/2-3/2 to base $3/2$, which will allow us to characterize the fixed points of $J_{3}$ in a much simpler way. Here, we do not refer to the relatively well-known fractional number system in base $3/2$ with digits 0 or $1$. To recall how such a system works, let us choose the natural number $4$ and represent it in base $3/2$. Clearly, a good default approximation of this number is $1000.01001001_{3/2}$ because
@@
where $d_{i}\in\mathcal{D}=\{0,1,2\}$ for some $k\in\mathbb{Z}_{+}$. If $k\geq 1$, we always assume that $d_{k}\neq 0$. When $N$ has the expansion of the form (8) we denote: $N=(d_{k}\cdots d_{0})_{3/2}.$ Note further that this expansion allows the digit $2$, which is not permitted in the standard base $3/2$ representation. Indeed, the expansion of the number $4$ in the form of (8) is $(212)_{3/2}$, since
@@
We now describe the process for computing an expansion of the form (8). To this end, we use the following four-step inductive algorithm, presented in [1, p. 61]:
Step 1: Set $N_{0}=N$ and write $2N_{0}=3N_{1}+d_{0}$ with $d_{0}\in\mathcal{D}$ and $N_{1}\in\mathbb{Z}_{+}$. Note that $d_{0}$ is the unique digit in $\mathcal{D}$ satisfying the congruence $d_{0}={\rm mod}(2N_{0},3)$ and thus, $N_{1}$ is also uniquely determined.
Step 2: Define the non-negative integers $N_{1},N_{2},\ldots$ recursively by the equations $2N_{i}=3N_{i+1}+d_{i},$ with $d_{i}\in\mathcal{D}$, uniquely defined as $d_{i}={\rm mod}(2N_{i},3)$.
Step 3: We now prove by induction on $i$ that
@@
Substituting this into the inductive hypothesis gives us:
@@
From the recurrence relation
$N_{i+1}=\frac{2}{3}N_{i}-\frac{1}{3}d_{i}\leq\frac{2}{3}N_{i},$
the sequence $\{N_{i}\}$ is strictly decreasing whenever $N_{i}\geq 3$. This guarantees that, after finitely many steps, we will reach a value $N_{i_{0}}\in\{1,2\}$ for some $i_{0}\geq 1$. We consider the two possible cases: (i) If $N_{i_{0}}=2$ then $N_{i_{0}+1}=1$ and $N_{i_{0}+2}=0$, so the expansion terminates at $\hat{i}=i_{0}+2$. (ii) If $N_{i_{0}}=1$ then $N_{i_{0}+1}=0$, hence, $\hat{i}=i_{0}+1$.
Therefore, for any initial value $N=N_{0}\in\mathbb{N}$, the sequence $\{N_{i}\}$ eventually reaches zero, which confirms that $N$ has a finite expansion of the form (8). Since each digit $d_{i}$ is uniquely determined by the congruence $d_{i}\equiv 2N_{i}\pmod{3}$ and must lie in $\mathcal{D}=\{0,1,2\}$, it follows that the expansion is unique.
The next table shows the $3/2$ expansions of the type (8) for the first twenty fixed points of $J_{3}$.
@@
Table 3: Values of $n_{p}^{(\ell)}$, $\overline{m}_{\ell}$, and their base $\frac{3}{2}$ expansions for $\ell=1,\ldots,20$.
@@ 2.3 Characterization of $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ through the base $3/2$ number system
Indeed, Table 3 exhibits several interesting features:
@@
    The base $3/2$ expansion of a fixed point $n_{p}^{(\ell)}$ repeats in the initial part of the base $3/2$ expansion of the next fixed point $n_{p}^{(\ell+1)}$.
@@
    The number of digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ is equal to the number of digits in the expansion of $n_{p}^{(\ell)}$ plus the number of pure high extremal points between them, $\overline{m}_{\ell}$, plus one.
@@
    The last $\overline{m}_{\ell}+1$ digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ are updated as follows:
@@
      The expansion is extended by the digit $1$ if $\overline{m}_{\ell}=0$.
@@
      The digits $02$ are appended to the expansion if $\overline{m}_{\ell}=1$.
@@
      The expansion is followed by the digit strings $012$, $0112$, $01112$, $\ldots$ if $\overline{m}_{\ell}=2,3,4,\ldots$.
@@
The total number of digits in the base $3/2$ expansion of the fixed point $n_{p}^{(\ell+1)}$ is equal to the number of digits in the base $3/2$ expansion of its predecessor, $n_{p}^{(\ell)}$, plus $\overline{m}_{\ell}+1$.
@@
To prove (i) let $\overline{m}_{\ell}=0$. Then, by (2),
@@
So, $d_{0}=1$, and in the induction formula (9), with $N=n_{p}^{(\ell+1)}$ and $N_{1}=n_{p}^{(\ell)}$ proving this case.
To prove (ii) and (iii), assume that $\overline{m}_{\ell}\geq 1$. We set $N_{0}=n_{p}^{(\ell+1)}.$ For simplicity, we define
@@
Therefore, $d_{1}={\rm mod}(2N_{1},3)=0$. Hence, $d_{1}=0$ and the digits 2 and 0 are added to the right of the base $3/2$ expansion of $N_{2}=n_{p}^{(\ell)}$ to represent $n_{p}^{(\ell+1)}$, establishing item (ii).
If $\overline{m}_{\ell}\geq 2$, we can repeat the algorithm $\overline{m}_{\ell}-1$ times starting at:
Step 1: Here, we have
@@
Until, Step $\overline{m}_{\ell}-1$: where we get,
$2N_{\overline{m}_{\ell}-1}=3A_{\overline{m}_{\ell}-1}-2=3(A_{\overline{m}_{\ell}-1}-1)+1.$
@@
Finally, by applying the algorithm over again, we get
$2N_{\overline{m}_{\ell}}=2A_{\overline{m}_{\ell}-1}-2=A_{\overline{m}_{\ell}}-2=3n_{p}^{(\ell)}.$
Thus,
$d_{\overline{m}_{\ell}}=0\quad\mbox{and}\quad N_{\overline{m}_{\ell}+1}=n_{p}^{(\ell)}$
proving item (iii).
The item (iv) follows directly from (i), (ii), and (iii), as each case increases the digit count of the base $3/2$ expansion by exactly $\overline{m}_{\ell}+1$. ∎
This result is a natural generalization of the explicit formula for the fixed points of the Josephus function $J_{2}$, discussed at the beginning of this section. Moreover, [6, 5] show that the binary expansion of $n$ can also be used to compute $J_{2}(n)$ via a digit shift. Specifically, if $n=(b_{k}\cdots b_{0})_{2}$ and $j:=\max\left\{j\in\{k-1,\ldots,0\right\}\mid b_{j}\neq 0\}$, then $J_{2}(n)=(b_{j}\cdots b_{0}b_{k})_{2}$. It remains open whether: (i) the base $3/2$ expansions of $n$ and $n_{p}^{(\ell+1)}$, together with the information of $\overline{m}_{\ell}$, can be used to solve the Josephus problem, $J_{3}(n)$. Equations (9) and (10) in [2] may provide a method to determine the base $3/2$ expansion of $J_{3}(n)$; (ii) the fixed points of the Josephus function $J_{4}$ admit a recursive digit formula when expressed in base $4/3$ number system. Exploring such a representation could reveal structural patterns analogous to those observed in the base $3/2$ expansions of the fixed points of $J_{3}$, which in turn generalize the case of $J_{2}$ in base $2$.

### v3

@@ Abstract
In this paper, we investigate properties of the fixed point sequence of the Josephus function $J_{3}$. First, we establish a connection between this sequence and the Chinese Remainder Theorem. Next, we identify a clear numerical pattern for the digits of two consecutive fixed points when they are written in a non-standard fractional number system in base $3/2$. This result enables us to derive a recursive procedure for determining the digits of their base $3/2$ expansions.
@@ 1 Introduction
This work investigates properties of the Josephus function $J_{3}$, which arises from the classical Josephus problem, a well-known combinatorial puzzle originating from a narrative by the Jewish-Roman historian Flavius Josephus (see [4, Book 3, Chapter 8, Part 7]). The problem is described as follows: Given a group of $n$ people seated at a round table, numbered consecutively clockwise from $1$ to $n$, begin counting from position $1$, skip two people, and eliminate the third. Continue this process with the remaining participants, always starting with the next person, skipping two, and removing the third, proceeding clockwise until only one person remains. This person is called the survivor.
A natural question is where one should initially sit to avoid elimination. Let $J_{3}(n)$ denote the position of the survivor. Computing $J_{3}(n)$ directly for large $n$ can be computationally expensive, so it is useful to identify properties of $J_{3}$ that can be exploited; see [2, 3].
@@
Definition 1 (High extremal points).
@@
Next, we outline the main results of this paper, organized into three subsections. We begin by motivating the main goal by recalling the well-known explicit formula for the fixed points of $J_{2}$. Subsection 2.1 connects the search for fixed points of $J_{3}$ with the Chinese Remainder Theorem by representing each fixed point as a solution of a system of congruences with pairwise coprime moduli. Subsection 2.2 introduces a fractional number system in base $3/2$, following [1, 7], and studies its relevance to fixed points. Finally, Subsection 2.3 uses this system to recursively determine the digits in the base $3/2$ expansion of each fixed point of $J_{3}$.
@@ 2 Main Results and Motivations
In this paper, we aim to derive a recursive formula for the digits of consecutive fixed points of $J_{3}$ in a suitable base. Such a formula was obtained in full for the special case of $J_{2}$ in [5, 6]. In this case, the fixed points are given by the explicit formula $n_{p}^{(\ell)}=2^{\ell}-1$, for all $\ell\in\mathbb{N}$. Their expansions in base $2$ exhibit a particularly simple recursive pattern. Specifically, if
@@ 2.1 Fixed points of $J_{3}$ and the Chinese Remainder Theorem
which tells us that $x_{1}=n_{p}^{(\ell+1)}$ and $x_{2}=n_{p}^{(\ell)}$ each satisfy the following system of linear congruences:
@@
where $p:=\overline{m}_{\ell}$ and $q:=\overline{m}_{\ell+1}$. Note that when $pq=0$, our argument does not apply. From this point on, we assume that $p,q\geq 1$. One can prove by induction on $p$ and $q$ that
@@
where $\gcd$ denotes the greatest common divisor. In this way, we uncover a connection between the sequence $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ and the CRT. Moreover, since the explicit expression for $z$ depends on a solution to Bézout’s identity, it is natural to examine how $1$ can be expressed as a linear combination of $3^{p}$ and $2^{q}$.
@@ 1 $1=3^{p}(1)+2^{1}\left(\tfrac{1}{2}(1-3^{p})\right),\quad\text{if }p\geq 1$
Table 2 provides Bézout’s linear combinations for all $p\in\mathbb{N}$ and $q=1,2,3,4,5$ in a compact form. We invite the interested reader to extend Table 2 to other values of $q$.
Remark 1.
@@
The first example corresponds to $n_{p}^{(17)}=3\,986\,218$ and $n_{p}^{(18)}=102\,162\,424$ given in [3, Table 1] or Table 3, where $p=7$ (i.e., $\overline{m}_{17}=7$) and $q=1$ (i.e., $\overline{m}_{18}=1$). According to (5) and Table 2, $a_{1}=1093$, $a_{2}=0$, and $x=1$, y=1093y=-1093. Therefore,
@@
Thus, $n_{p}^{(18)}=102\,162\,424$ must be congruent to $3280$ modulo $4374$, which is indeed the case since (102 162 4243280)/4374=23356(102\,162\,424-3280)/4374=23356.
The second example corresponds to $n_{p}^{(13)}=46084$ and $n_{p}^{(14)}=103690$; see Table 3 below. In this case, $p=1$ (i.e., $\overline{m}_{13}=1$) and $q=5$ (i.e., $\overline{m}_{14}=5$). Here, $a_{1}=1$, $a_{2}=10$ in (5), and $x=11$, y=1y=-1 for Bézout’s identity (6) since
@@ 2.2 Expansions in base $3/2$
which provides motivation for studying representations of natural numbers in base $3/2$. Inspired by (7) and the basic conceptual framework presented in the introduction of [7, Section 3], we move from base 3/2-3/2 to base $3/2$, which will allow us to characterize the fixed points of $J_{3}$ in a much simpler way. Here, we do not refer to the well-known fractional number system in base $3/2$ with digits 0 or $1$. To illustrate how such a system works, let us represent the natural number $4$ in base $3/2$. A natural approximation is $1000.01001001_{3/2}$, since
@@
where $d_{i}\in\mathcal{D}=\{0,1,2\}$ for some $k\in\mathbb{Z}_{+}$. If $k\geq 1$, we always assume that $d_{k}\neq 0$. When $N$ has an expansion of the form (8), we denote $N=(d_{k}\cdots d_{0})_{3/2}.$ Note that this expansion allows the digit $2$, which is not permitted in the standard base $3/2$ representation. Indeed, the expansion of the number $4$ in the form of (8) is $(212)_{3/2}$, since
@@
We now describe the process for computing an expansion of the form (8) via the following four-step algorithm, presented in [1, p. 61]:
Step 1: Set $N_{0}=N$ and write $2N_{0}=3N_{1}+d_{0}$ with $d_{0}\in\mathcal{D}$ and $N_{1}\in\mathbb{Z}_{+}$. Note that $d_{0}$ is the unique digit in $\mathcal{D}$ satisfying $d_{0}\equiv 2N_{0}\pmod{3}$, and thus $N_{1}$ is also uniquely determined.
Step 2: Define the non-negative integers $N_{1},N_{2},\ldots$ recursively by $2N_{i}=3N_{i+1}+d_{i}$, with $d_{i}\in\mathcal{D}$ uniquely defined by $d_{i}\equiv 2N_{i}\pmod{3}$.
Step 3: We prove by induction on $i$ that
@@
Substituting this into the inductive hypothesis gives:
@@
From the recurrence relation $N_{i+1}=\frac{2}{3}N_{i}-\frac{1}{3}d_{i}\leq\frac{2}{3}N_{i}$, the sequence $\{N_{i}\}$ is strictly decreasing whenever $N_{i}\geq 3$. This guarantees that, after finitely many steps, we reach a value $N_{i_{0}}\in\{1,2\}$ for some $i_{0}\geq 1$. We consider two cases: (i) if $N_{i_{0}}=2$, then $N_{i_{0}+1}=1$ and $N_{i_{0}+2}=0$, so the expansion terminates at $\hat{i}=i_{0}+2$; (ii) if $N_{i_{0}}=1$, then $N_{i_{0}+1}=0$, hence $\hat{i}=i_{0}+1$.
Therefore, for any $N=N_{0}\in\mathbb{N}$, the sequence $\{N_{i}\}$ eventually reaches zero, confirming that $N$ has a finite expansion of the form (8). Since each digit $d_{i}$ is uniquely determined by $d_{i}\equiv 2N_{i}\pmod{3}$ and must lie in $\mathcal{D}=\{0,1,2\}$, the expansion is unique.
The following table displays the base $3/2$ expansions of the form (8) for the first twenty fixed points of $J_{3}$.
@@
Table 3: Values of $n_{p}^{(\ell)}$, $\overline{m}_{\ell}$, and the corresponding base $3/2$ expansions for $\ell=1,\ldots,20$.
@@ 2.3 Characterization of $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ through the base $3/2$ number system
Indeed, Table 3 exhibits several notable features:
@@
    The base $3/2$ expansion of a fixed point $n_{p}^{(\ell)}$ appears as the initial segment of the base $3/2$ expansion of the next fixed point $n_{p}^{(\ell+1)}$.
@@
    The number of digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ equals the number of digits in the expansion of $n_{p}^{(\ell)}$ plus $\overline{m}_{\ell}+1$, where $\overline{m}_{\ell}$ is the number of pure high extremal points between them.
@@
    The last $\overline{m}_{\ell}+1$ digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ are determined as follows:
@@
      The digit $1$ is appended if $\overline{m}_{\ell}=0$.
@@
      The digits $02$ are appended if $\overline{m}_{\ell}=1$.
@@
      The digit strings $012$, $0112$, $01112$, $\ldots$ are appended if $\overline{m}_{\ell}=2,3,4,\ldots$, respectively.
@@
The total number of digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ equals the number of digits in the base $3/2$ expansion of $n_{p}^{(\ell)}$ plus $\overline{m}_{\ell}+1$.
@@
To prove (i), let $\overline{m}_{\ell}=0$. Then, by (2),
@@
So $d_{0}=1$, and in the induction formula (9) with $N=n_{p}^{(\ell+1)}$ and $N_{1}=n_{p}^{(\ell)}$, this proves (i).
To prove (ii) and (iii), assume that $\overline{m}_{\ell}\geq 1$. Set $N_{0}=n_{p}^{(\ell+1)}$ and define
@@
Therefore, $d_{1}={\rm mod}(2N_{1},3)=0$, and the digits $2$ and 0 are appended to the right of the base $3/2$ expansion of $N_{2}=n_{p}^{(\ell)}$ to represent $n_{p}^{(\ell+1)}$, establishing (ii).
If $\overline{m}_{\ell}\geq 2$, we repeat the algorithm $\overline{m}_{\ell}-1$ times starting at:
Step 1: We have
@@
Continuing in this way, at Step $\overline{m}_{\ell}-1$ we obtain
$2N_{\overline{m}_{\ell}-1}=3A_{\overline{m}_{\ell}-1}-2=3(A_{\overline{m}_{\ell}-1}-1)+1,$
@@
Finally, applying the algorithm once more gives
$2N_{\overline{m}_{\ell}}=2A_{\overline{m}_{\ell}-1}-2=A_{\overline{m}_{\ell}}-2=3n_{p}^{(\ell)},$
so
$d_{\overline{m}_{\ell}}=0\quad\mbox{and}\quad N_{\overline{m}_{\ell}+1}=n_{p}^{(\ell)},$
proving (iii).
Item (iv) follows directly from (i), (ii), and (iii), since each case increases the digit count by exactly $\overline{m}_{\ell}+1$. ∎
This result is a natural generalization of the explicit formula for the fixed points of $J_{2}$, discussed at the beginning of this section. Moreover, [5, 6] show that the binary expansion of $n$ can be used to compute $J_{2}(n)$ via a digit shift. Specifically, if $n=(b_{k}\cdots b_{0})_{2}$ and $j:=\max\left\{j\in\{k-1,\ldots,0\}\mid b_{j}\neq 0\right\}$, then $J_{2}(n)=(b_{j}\cdots b_{0}b_{k})_{2}$.
Two directions remain open. First, whether the base $3/2$ expansions of $n$ and $n_{p}^{(\ell+1)}$, together with $\overline{m}_{\ell}$, can be used to compute $J_{3}(n)$ directly; equations (9) and (10) in [2] may provide a starting point. Second, whether the fixed points of $J_{4}$ admit a recursive digit formula in base $4/3$; such a representation could reveal structural patterns analogous to those found for $J_{3}$ in base $3/2$, which in turn generalize the case of $J_{2}$ in base $2$.